# Distinctive Feature Codec: Adaptive Segmentation for Efficient Speech Representation

### v1

@@ D.1 Details of Semantic Information Distillation
Table 4: Impact of Semantic Distillation (SD) on speech tokenization models. SD refers to the process of guiding the first RVQ layer with HuBERT representations. Lower values are better for MEL Error, STFT, and WER; higher values are better for PSEQ and STOI.
@@
To investigate the impact of semantic distillation on speech tokenization models, we conducted experiments comparing performance with and without distillation across different frame rates. Table 4 presents these results, revealing several important insights not previously reported in the original SpeechTokenizer work.
Our findings demonstrate a consistent pattern: when semantic distillation is applied, reconstruction quality metrics (MEL Error, STFT, PSEQ, STOI) tend to decrease, while speech content preservation measured by WER improves. This trade-off is evident in both our Distinctive Codec at 9.5 Hz and SpeechTokenizer at 50 Hz, where the addition of semantic distillation increases reconstruction errors but significantly reduces word error rates.
@@
Appendix E Many-to-one Encoding and Decoding with Distinctive Feature Autoencoding
In our proposed framework, we introduce a many-to-one quantization strategy that processes variable-length segments based on distinctive features. This strategy is implemented via two key procedures—encoding and decoding—which are tightly integrated with our distinctive feature segmentation and the PerSegmentAutoEncoder. The following subsections detail these procedures.
E.1 Encoding Procedure
Given a speech signal $\mathbf{x}\in\mathbb{R}^{B\times 1\times T}$, the encoding procedure comprises four main steps:
  1. 1.
    Feature Extraction and Dimensionality Reduction: The input signal is first processed by a SEANet encoder, yielding a latent representation:
    $\mathbf{e}=E_{\theta}(\mathbf{x})\in\mathbb{R}^{D\times T}.$ (19)
    The resulting feature is reshaped and projected via a linear layer (denoted as feature_decrease) to obtain a lower-dimensional embedding suitable for quantization:
    $\tilde{\mathbf{e}}\in\mathbb{R}^{B\times\text{quantizer\_dim}\times T}.$ (20)
  2. 2.
    Segmentation via Boundary Detection: A boundary detector identifies salient temporal boundaries, partitioning $\tilde{\mathbf{e}}$ into variable-length segments. For each sample $b$, let $\{b_{1}^{(b)},b_{2}^{(b)},\ldots,b_{M}^{(b)}\}$ denote the detected boundaries; using the convention $b_{0}^{(b)}=0$ and $b_{M+1}^{(b)}=T$, each segment is defined as:
    $\mathbf{s}_{i}^{(b)}=\tilde{\mathbf{e}}^{(b)}(:,\,b_{i-1}^{(b)}:b_{i}^{(b)}).$ (21)
  3. 3.
    Segment Compression via Autoencoding: Each segment $\mathbf{s}_{i}^{(b)}\in\mathbb{R}^{\text{quantizer\_dim}\times l_{i}}$, where $l_{i}$ denotes the segment length, is compressed using the PerSegmentAutoEncoder. This yields a fixed-length representation:
    $\mathbf{z}_{i}^{(b)}=f_{\text{AE}}^{\text{enc}}(\mathbf{s}_{i}^{(b)})\in\mathbb{R}^{H\times 1},$ (22)
    with $H$ representing the hidden dimension of the autoencoder.
  4. 4.
    Group-wise Quantization and Composite Token Construction: The compressed representation $\mathbf{z}_{i}^{(b)}$ is fed to a many-to-one FSQ quantizer. In our implementation, the high-dimensional vector is divided into $N$ groups, with each group quantized independently. Denote the discrete index for group $g$ as:
    $q_{i}^{(b),g}\in\{0,1,\ldots,L-1\},$ (23)
    where $L$ is the number of quantization levels (e.g., $L=8$). The composite token for the segment is then computed via a positional expansion:
    $c_{i}^{(b)}=\sum_{g=0}^{N-1}q_{i}^{(b),g}\cdot L^{g}.$ (24)
    The scalar $c_{i}^{(b)}$ uniquely represents the segment in a discrete vocabulary of size $L^{N}$.
The encoding procedure returns, for each sample, a list of composite tokens corresponding to its segments along with the associated segmentation metadata (i.e., segment start, end, and length).
E.2 Decoding Procedure
Given the composite tokens and segmentation information obtained from the encoding stage, the decoding procedure reconstructs the original speech signal by reversing the encoding steps:
  1. 1.
    Composite Token Decomposition: For each composite token $c$ corresponding to a segment, we recover the $N$ discrete indices using modulus and integer division:
    $q^{g}=\left\lfloor\frac{c}{L^{g}}\right\rfloor\mod L,\quad g=0,1,\ldots,N-1.$ (25)
    Each recovered $q^{g}$ is reconstituted as a tensor suitable for the inverse mapping of the FSQ module.
  2. 2.
    Reconstruction of Compressed Representations: Each index $q^{g}$ is mapped back to a continuous latent code via the FSQ module’s inverse function and an associated expansion layer. The continuous latent codes from all groups are then concatenated to form the reconstructed compressed representation:
    $\hat{\mathbf{z}}_{i}^{(b)}\in\mathbb{R}^{H\times 1}.$ (26)
  3. 3.
    Segment Reconstruction via Autoencoding: The reconstructed compressed representation is then passed through the PerSegmentAutoEncoder decoder to recover the original segment:
    $\hat{\mathbf{s}}_{i}^{(b)}=f_{\text{AE}}^{\text{dec}}(\hat{\mathbf{z}}_{i}^{(b)},\,l_{i})\in\mathbb{R}^{\text{quantizer\_dim}\times l_{i}},$ (27)
    where $l_{i}$ is the original length of the segment.
  4. 4.
    Temporal Assembly and Final Reconstruction: The reconstructed segments $\{\hat{\mathbf{s}}_{i}^{(b)}\}$ are temporally reassembled according to their segmentation metadata to form the complete latent representation:
    $\hat{\mathbf{e}}^{(b)}\in\mathbb{R}^{\text{quantizer\_dim}\times T}.$ (28)
    Finally, a linear projection (denoted as feature_increase) is applied, followed by the SEANet decoder, to produce the final reconstructed waveform:
    $\hat{\mathbf{x}}^{(b)}=D_{\phi}(\hat{\mathbf{e}}^{(b)}).$ (29)
Equations 24 and 25 encapsulate the critical operations ensuring a bijective mapping between the per-group quantization indices and the composite token. This guarantees that the decoding process can accurately recover the compressed representations, thereby enabling faithful reconstruction of the original speech signal.
Appendix F Pseudocode of Group-wise Scalar Quantization (GSQ)
Algorithm 1 Many2One Group-wise Scalar Quantization (Many2oneFsq)
0:  Input tensor $x\in\mathbb{R}^{B\times T\times D}$, group count $G$, group size $d=D/G$
1:  for $i=1$ to $G$ do
2:     $x_{i}\leftarrow x[:,:,i\cdot d:(i+1)\cdot d]$Split input into $G$ groups}
3:     $c_{i}\leftarrow W_{i}^{(c)}x_{i}$Compress to scalar via linear layer}
4:     $\hat{c}_{i},\text{index}_{i}\leftarrow\text{FSQ}_{i}(c_{i})$Quantize scalar}
5:     $\hat{x}_{i}\leftarrow W_{i}^{(e)}\hat{c}_{i}$Expand back to group dimension}
6:  end for
7:  $\hat{x}\leftarrow\text{Concat}([\hat{x}_{1},\ldots,\hat{x}_{G}],\text{dim}=-1)$
8:  return  $\hat{x}$, $\{\text{index}_{i}\}_{i=1}^{G}$
Algorithm 2 Many2Many Group-wise Scalar Quantization (Many2manyFsq)
0:  Input tensor $x\in\mathbb{R}^{\cdot\times D}$, chunk size $d$, number of chunks $N=D/d$
1:  $x\leftarrow\text{Split}(x,\text{chunk\_size}=d)$Split input into $N$ sub-vectors}
2:  for $i=1$ to $N$ do
3:     $q_{i},\text{index}_{i}\leftarrow\text{FSQ}_{i}(x_{i})$Apply FSQ to each chunk}
4:  end for
5:  $q\leftarrow\text{Concat}([q_{1},\ldots,q_{N}],\text{dim}=-1)$
6:  $\text{indices}\leftarrow\text{Stack}([\text{index}_{1},\ldots,\text{index}_{N}],\text{dim}=-1)$
7:  return  $q$, indices
Appendix G Clarification on Distinctive Boundary Detection
One potential concern from readers may stem from the role of boundary detection in our framework. Since our method relies on identifying perceptually significant boundaries to construct variable-length segments, it may appear that our model depends heavily on the quality or supervision of these boundaries.
However, a key motivation behind our use of distinctive features is precisely to reduce the reliance on strong linguistic priors (e.g., phoneme or word boundaries), which are often language-specific and costly to annotate. Our boundary detector is trained in a fully self-supervised manner using acoustic contrastive learning, which enables the model to discover transition points purely based on local signal variations—without requiring any phoneme-level or syllable-level supervision.
This perceptually driven approach offers two main advantages: (1) it generalizes well across different languages and speaking styles, and (2) it aligns naturally with the non-uniform information distribution in speech. While we do not provide a standalone metric for boundary detection accuracy due to the absence of human-annotated acoustic transitions, our experiments—including the ablation study and codebook utilization analysis—demonstrate that the learned boundaries improve both compression efficiency and representation quality.
We view this property not as a limitation, but as a strength of our design: by avoiding handcrafted linguistic rules, the boundary detection process becomes more scalable, adaptable, and compatible with large-scale unlabeled speech data.
Appendix H Generalization to Out-of-Domain and Code-Switched Speech
To evaluate the generalization capability of our model beyond English and the LibriSpeech training domain, we conduct a zero-shot inference experiment on the SEAME dataset [69], a 200-hour spontaneous Mandarin-English code-switching corpus widely used in multilingual speech research. SEAME poses three primary challenges,1.Conversational spontaneity: frequent disfluencies (hesitations, overlaps, false starts) and colloquial prosody.2.Telephony distortions and environmental noise: low-bitrate channel artifacts and background sounds typical of mobile or landline settings.3.Cross-lingual code-switching: rapid alternation between Mandarin and English within utterances
Without any fine-tuning, we directly apply our Distinctive Codec model trained on LibriSpeech to a subset of 50 utterances randomly selected from SEAME. The evaluation focuses on PSEQ, which serves as the primary perceptual quality metric and correlates strongly with both intelligibility and signal fidelity.
Under a low token rate setting (9.5 Hz), our model achieves a PSEQ score of 1.4214, significantly outperforming SpeechTokenizer’s PSEQ score of 1.0695 under the same configuration. This substantial improvement demonstrates that the proposed distinctive feature-based tokenization not only preserves critical acoustic cues in the English domain but also exhibits robust transferability to unseen languages and mixed-lingual acoustic conditions.
Appendix I Comparison with WaveTokenizer
To further validate the effectiveness of our distinctive feature-based approach, we conducted additional experiments comparing our method with WaveTokenizer [70], another state-of-the-art neural speech codec. While our main results (Table 1) demonstrate competitive performance against SpeechTokenizer, this additional comparison provides broader context for our approach within the current landscape of speech tokenization methods.
For a fair comparison, we trained WaveTokenizer on the LibriSpeech dataset with a token rate of 10 Hz, matching the operating conditions of our Distinctive Codec (9.5 Hz). The results are presented in Table 5.
Table 5: Performance comparison between WaveTokenizer and our Distinctive Codec at similar token rates. Lower values are better for MEL Error, STFT, and WER; higher values are better for PESQ and STOI.
Model Token Rate Metrics
[…]
WaveTokenizer 10 Hz 0.3139 1.0459 1.8333 0.7449 0.5535
Distinctive Codec (GSQ) 9.5 Hz 0.2006 0.6021 2.1901 0.8114 0.3061

### v2

@@ VI-E Investigate the impact of semantic distillation
TABLE VI: Impact of Semantic Distillation (SD) on speech tokenization models. SD refers to the process of guiding the first RVQ layer with HuBERT representations. Lower values are better for MEL Error, STFT, and WER; higher values are better for PESQ and STOI.
@@
To investigate the impact of semantic distillation on speech tokenization models, we conducted experiments comparing performance with and without distillation across different frame rates. Table VI presents these results, revealing several important insights not previously reported in the original SpeechTokenizer work.
Our findings demonstrate a consistent pattern: when semantic distillation is applied, reconstruction quality metrics (MEL Error, STFT, PESQ, STOI) tend to decrease, while speech content preservation measured by WER improves. This trade-off is evident in both our Distinctive Codec at 9.5 Hz and SpeechTokenizer at 50 Hz, where the addition of semantic distillation increases reconstruction errors but significantly reduces word error rates.
@@
VII Conclusion
In this paper, we investigated the effectiveness of distinctive features for speech representation in depression detection, demonstrating that adaptive segmentation aligned with acoustic boundaries preserves critical temporal dynamics. Our experiments address the fundamental limitation of fixed-rate processing that destroys timing information essential for clinical applications. Results show that distinctive feature-based tokenization produces more coherent latent representations with improved codebook utilization, while our Group-wise Scalar Quantization strategy enables stable quantization at low token rates. The substantial performance gains in depression detection (+35.0% relative improvement) validate that preserving temporal structure is crucial for capturing clinically relevant biomarkers such as pause patterns and speech rate variations. This work establishes distinctive features as a promising direction for neural speech codecs in temporally sensitive applications, with potential benefits for clinical assessment systems and other domains requiring faithful preservation of timing dynamics.
Acknowledgement
This work was supported by the Australian Research Council Discovery Project DP230101184.